The length of a rectangular playing field is 5 feet less than it's width. If the perimeter of the playing field is 230 feet, find the length and width of the field.
Answer by elima(1433)

P=2(L+W)
P=230
L=w-5
230=2(w-5+w)
230=2(2w-5)
230=4w-10
240=4w
60=w
L=60-5=55
:)